OpenStudy (anonymous):

whats the answer to 0.5x-1.7=0.3x-1

OpenStudy (anonymous):

.5x-.3x=1.7-1

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

arent you understanding it yet

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

k

OpenStudy (anonymous):

0.2x=0.7 x=0.7/0.2=7/2=3.5

OpenStudy (anonymous):

Q: what is the answer to 0.2(7+2t)=0.4t+1.4 That answer is -17.5 and here is why 0.2(7+2t)=0.4t+1.4 7+2t = 0.4t+1.4/0.2 now 2t-0.4t = 1.4-7/0.2 now 1.6t = -5.6/0.2 now 1.6t = -28 now t = -28/1.6 now t = -17.5